OpenStudy (anonymous):

A gem store sells beads made of amber and quartz. For 4 amber beads and 4 quartz beads, the cost is $46.00. For 1 amber bead and 3 quartz beads, the cost is $14.50. Find the price of each type of bead.

OpenStudy (anonymous):

if you put say \(x\) as the price of the amber bead and \(y\) as the price of the quartz bead you get the two equations \[4x+4y=46\\x+3y=14.50\]

OpenStudy (anonymous):

the first equation you can divide all by 2 and get \[2x+2y=23\] and the second you can solve for \(x\) and get \(x=14.5-3y\)

OpenStudy (anonymous):

i will shut up and let you finish

OpenStudy (anonymous):

or else just add that if you substitute \(x\) in the first equation by \(14.5-3y\) you get \[2(14.5-3y)+2y=23\] and you can solve that for \(y\)

OpenStudy (anonymous):

Is y=1.5 ?

OpenStudy (aaronq):

yeah

OpenStudy (anonymous):

Thanks guys!

OpenStudy (aaronq):

np